OpenStudy (anonymous):

A set of equations is given below: Equation A: y = x + 1 Equation B: y = 4x + 5 Which of the following steps can be used to find the solution to the set of equations? x + 1 = 4x + 5 x = 4x + 5 x + 1 = 4x x + 5 = 4x + 1

OpenStudy (anonymous):

@asdfghjkl24562

OpenStudy (michele_laino):

here we can apply the elimination method, so we can write: \[\Large \begin{gathered} x + 1 = y = 4x + 5 \hfill \\ \hfill \\ x + 1 = 4x + 5 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

I still don't don't get it I mean I know I got to subtract from both sides

OpenStudy (michele_laino):

no, please, we have these two equations: \[\Large\begin{gathered} y = 4x + 5 \hfill \\ y = x + 1 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

now the left sides are equal each other

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

then also the right sides have to be equal each other

OpenStudy (michele_laino):

so we can write: \[\Large x + 1 = 4x + 5\]

OpenStudy (anonymous):

can you help me with 2 more test there only 5 questions each and ill give you meadls

OpenStudy (michele_laino):

another reasoning is: \[x + 1 = y\] and \[y = 4x + 5\] then if we apply the transitive property of equality, we get: \[x + 1 = 4x + 5\]

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

I can help you!